Ticket Number: HEA-NOC/20070209-5 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: DCU
Ticket Opened: 20070209 15:28 UTC Problem Start: 20070209 13:00 UTC
Ticket Closed: 20070227 09:48 UTC Problem End: 20070226 09:48 UTC
Site/line: Crumlin VEC, DCU, Hermes, IT Blanchardstown, Kilester VEC,
Liberties VEC, NQAI, Schools Network, TCD, Whitehall VEC,
BH-DCU, Bluenet CW-DCU, Bluenet TCD-DCU, DCU-POP PoP
DCU have advised that there will be planned electrical work carried out in
the location of the HEAnet PoP on the weekend of the 24th - 25th of
This work should be expected to cause a complete outage to all equipment
and circuits terminated in the DCU PoP while being carried out on the 24th
or 25th of February. It has been estimated that this should take 30 - 60
There will be no connectivity to the following customers for the duration
of the work:
DCU, Hermes, NQAI, Whitehall, Killester, Crumlin, Liberties.
Redundancy to the following customers will be affected:
TCD, IT Blanchardstown, HEAnet Schools (Blanchardstown), HEAnet Services
20070209 15:04 BS opening advisory ticket. This ticket will be updated with
more information once received.
20070223 15:04 BS outage will be taking place between 8:00am and 9:00am
20070224 09:10 EK Power was returned tothe HEAnet PoP at 9:04. The outage
was from 8:13 to 9:04. All clients connected to the DCU PoP are now back
up except for DCU.DCU is expected back up around 17:30 today.
2 out of the 3 Secondary DNS servers are resolving DCU address to the
correct temporary HEAnet website. So 1 in 3 requests to DCU websites will
fail to reach the temporary maintenance page. This will be resolved after
20070224 17:52 EK DCU is reachable since 14:55. The majority of servers
and services at DCU are now back up. Some servers/services will remain
unavailable until tomorrow.
20070227 09:48 BS closing ticket, work completed.
This ticket can be monitored at http://www.hea.net/tickets/20070209-5